OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 2y = 12 x – y = 3

hero (hero):

4x + 2y = 12 x - y = 3 y = (12 - 4x)/2 y = x - 3 y = y 2(x-3) = 12 - 4x 2x - 6 = 12 - 4x 4x + 2x = 12 + 6 6x = 18 x = 18/6 x = 3

hero (hero):

The advantage of isolating is y first is so that you can check your answer by graphing as well.